CS 582 — Database Management System II — Spring 2007

Quiz 2

NAME:

Given the following database schema about computer manufacturers and their products.

• Product(maker, \underline{model} , type)
• PC(\underline{model} , speed, ram, hd, cd, price)
• Laptop(\underline{model} , speed, ram, hd, screen, price)

We assume that model number for each type of computer/laptiop is unique and hence can be used as a key for every relations. In what follows, a computer is understood as either a laptop or a PC. Write relational algebra expression for the following queries:

• List all computer models whose speed is at least 288 (or 2.88MHz).

• List all makers (or manufactures) who produce at least two models.